OpenStudy (anonymous):

Solve |2y - 2| = y. A.{2/3, 2} B.{2, 3} C.{2, 5} D.{1/3, 2}

OpenStudy (anonymous):

d

OpenStudy (anonymous):

1) \(2y-2=y \implies y=2 \). 2) \(2y-2=-y \implies 3y=2 \implies y=\frac{2}{3}\).

OpenStudy (anonymous):

A is the answer.